The time it takes to assemble a car in a certain factory is a random variable with a mean of 20 hours and a standard deviation of 2 hours.
If 25 cars are assembled per day:
1. How many cars are assembled in less than 24 hours?
2. How many cars are assembled in atleast 22 hours?
1
Expert's answer
2021-10-12T05:07:21-0400

Let "X=" the time it takes to assemble a car in a certain factory: "X\sim N(\mu, \sigma^2)."  

Given "\mu=20\ h, \sigma=2\ h."


"P(X<24)=P(Z<\dfrac{24-20}{2})=P(Z<2)"

"\approx0.977250"


"P(X\geq22)=1-P(X<22)="

"1-P(Z<\dfrac{22-20}{2})=1-P(Z<1)"

"\approx0.158655"

1.

"0.977250(25)=24(cars)"

2.


"0.158655(25)=4(cars)"